OpenStudy (anonymous):

Factorise completely: 1) 2mh- 2nh- 3mk+ 3nk Thank you!

OpenStudy (amoodarya):

2mh- 2nh- 3mk+ 3nk ( 2mh- 2nh) +(- 3mk+ 3nk ) =2h(m-n) -3k(m-n) now you can factor (m-n) can yo go on ?

OpenStudy (anonymous):

hmmm how to do that?

OpenStudy (amoodarya):

=2h(m-n) -3k(m-n) =(m-n)(2h-3k)

OpenStudy (anonymous):

thanks alot! i now understand!

OpenStudy (amoodarya):

I hope so your welcome